OpenStudy (anonymous):

2x squared - 5x =3 solve using the quadratic formula please

OpenStudy (anonymous):

@AccessDenied @Jadeishere @whpalmer4

OpenStudy (jadeishere):

so a = 2 b = -5 and c = -3

hero (hero):

\(2x^2 - 5x = 3\) \(2x^2 - 5x - 3 = 0\) 2 = a -5 = b -3 = c Plug those values into the quadratic formula: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (anonymous):

i don't know how to do the quadratic formula

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[2x ^{2}\]

hero (hero):

\[x = \frac{-(-5) \pm \sqrt{(-5)^2 - 4(2)(-3)}}{2(2)}\]

OpenStudy (anonymous):

what you do from there?

hero (hero):

Simplify it

OpenStudy (jadeishere):

so 2 = a, -5 = b and - 3 = c so -(-5) = 5 (-5)^2 - 4(2)(-3) = 25 + 24 = 49 so x = 5 +- the square root 0f 49 all over 2 times 2

OpenStudy (jadeishere):

\[x = \frac{ 5 \pm \sqrt {49} }{ 4} \]

OpenStudy (jadeishere):

the square root of 49 is 7 so x = 5 plus or minus 7 all over 4 x = 3 or x = -1/2

OpenStudy (anonymous):

thank you a million again

OpenStudy (jadeishere):

@Hero helped too

OpenStudy (anonymous):

true THANKS HERO YOUR MY HERO @Hero